OpenStudy (twalt13):

Can anyone help me ? I got stuck Find the derivative of f(x) = -12x^2 + 9x at x = 6.

OpenStudy (twalt13):

find f(x) = -12x^2+9x find f(x+h) = -12(x+h)^2 + 9(x+h) = -12x^2-12h^2+9x+9h

OpenStudy (twalt13):

Now i gotta put that into this formula\[f(x+h)-f(x)\div h\]

OpenStudy (twalt13):

so... -12x^2-12h^2+9x+9h -(-12x^2)+9x) -12x^2-12h^2+9x+9h +12x^2-9x combine like terms -12h+9h/h here is where i got stuck

OpenStudy (twalt13):

*-12h^2+9h/h

hartnn (hartnn):

find f(x+h) = -12(x+h)^2 + 9(x+h) = -12x^2-12h^2+9x+9h ********incorrect \((a+b)^2= a^2+2ab+b^2 \) where did your '2ab' term go ?

OpenStudy (twalt13):

idk i didnt know there was a 2ab term

hartnn (hartnn):

now you know, so whats (x+h)^2= ... ?

OpenStudy (twalt13):

@hartnn i solved that

hartnn (hartnn):

i am glad :)